OpenStudy (anonymous):

What is the solution to the rational equation 2x/3 - 4/3 = 6 ? thanks

OpenStudy (anonymous):

We must get x by itself. To do this, add \[\frac{4}{3}\] to both sides first.

OpenStudy (anonymous):

Then we have \[6+\frac{4}{3}=\frac{2}{3}x\]

OpenStudy (anonymous):

To get rid of the (2/3), simply multiply it by (3/2), because (2/3)*(3/2) = 1, and if we have that then we have x by itself. But remember, what you do to one side you must do to the other. So we have now: \[(6+\frac{4}{3})\frac{3}{2}=\frac{2}{3}\frac{3}{2}x=x\]

OpenStudy (anonymous):

So do the computation on the left, and that will be your final result.

OpenStudy (anonymous):

x5/y ?

OpenStudy (anonymous):

Typo

OpenStudy (anonymous):

(6+ (4/3))*(3/2) = 11

OpenStudy (anonymous):

=x

OpenStudy (anonymous):

ohh ok i see!

OpenStudy (anonymous):

Yeah sorry about that lol

OpenStudy (anonymous):

sorry about that i was looking at the wrong problem

OpenStudy (anonymous):

thanks again for going step by step :]